# Evaluation Framework for Highlight Explanations of Context Utilisation in Language Models

### v1

@@ Abstract
Context utilisation, the ability of Language Models (LMs) to incorporate relevant information from the provided context when generating responses, remains largely opaque to users, who cannot determine whether models draw from parametric memory or provided context, nor identify which specific context pieces inform the response. Highlight explanations (HEs) offer a natural solution as they can point the exact context pieces and tokens that influenced model outputs. However, no existing work evaluates their effectiveness in accurately explaining context utilisation. We address this gap by introducing the first gold standard HE evaluation framework for context attribution, using controlled test cases with known ground-truth context usage, which avoids the limitations of existing indirect proxy evaluations. To demonstrate the framework’s broad applicability, we evaluate four HE methods – three established techniques and MechLight, a mechanistic interpretability approach we adapt for this task – across four context scenarios, four datasets, and five LMs. Overall, we find that MechLight performs best across all context scenarios. However, all methods struggle with longer contexts and exhibit positional biases, pointing to fundamental challenges in explanation accuracy that require new approaches to deliver reliable context utilisation explanations at scale. 111github.com/lianyiyi/Transparent-Context-Usage 222huggingface.co/datasets/copenlu/transparent-context-usage
Equal contribution.
@@ 1 Introduction
Figure 1: Utility evaluation of two HEs in our framework under the Double-Conflicting context setup. In this example, the model selects the answer from passage two. Explainer 1 shows better utility than Explainer 2.
Language Models (LMs) operating on context-dependent tasks (QA, summarisation, dialogue modeling) lack transparency regarding whether  (Jin et al., 2024; Yu et al., 2023; Monea et al., 2024a) or how the provided context informed their generation. Highlight explanations (HEs) address this need naturally by pinpointing portions of the context responsible for the generation. See Fig.1 for an example of HEs with high/low accuracy. Although HEs have proven valuable for understanding model decisions across various tasks (Sun et al., 2025; Ray Choudhury et al., 2023; Atanasova et al., 2020), no existing work evaluates their effectiveness in accurately explaining context utilisation.
@@
Across five LMs and four commonly used context-usage datasets, we find that MechLight HEs perform best across all context scenarios. However, two systematic limitations persist across all HEs: (i) length sensitivity – HE accuracy degrades as context grows; and (ii) position biases under dual‑context inputs: FA/IG tend to favour later (near‑question) pieces, while ATTN/MechLight favour earlier pieces. Surprisingly, the widely used IG and MechLight exhibit poor accuracy in most context scenarios, rendering them useless in revealing the model’s context utilisation. These failures also underscore the urgent need for explanation techniques that maintain accuracy at scale and overcome positional biases in multi-document settings.
@@ 2.1 Studies of Context Usage
Upervised Context Usage Explanations. To attribute the model’s answer to the specific part of the context, SelfCite trains a classifier on pseudo‑citations generated by the LLM itself (Chuang et al., ); ContextCite scores each sentence by the drop in answer likelihood when it is masked (Cohen-Wang et al., 2024). They require extra supervision, expensive perturbations, and only explain at the sentence level.
@@ 3.3 Metrics
$\small\begin{split}\Delta Rank@k^{\mathrm{grp}}(T;A,B)&=Rank@k(T,D_{B})\\ &\qquad-Rank@k(T,D_{A})\end{split}$ (1)
@@
$\small\begin{split}\Delta Rank@k^{\mathrm{inst}}_{D_{C_{a}}}&=\frac{1}{|{D}_{C_{a}}|}\sum_{x\in{D}_{C_{a}}}(Rank@k^{\mathrm{inst}}(c_{b},x)\\ &-Rank@k^{\mathrm{inst}}(c_{a},x))\end{split}$ (2)
@@
$\small MDL-Bits@k=L_{\mathrm{preq}}(Y\mid X^{(k)}_{s})$ (4)
@@
$\small\mathrm{RR}(x)=\nicefrac{{1}}{{\mathrm{rank}\!\big(\text{Ans}_{\cdot};x\big)}}$ (5)
$\small MRR\big(T{=}\text{Ans}_{\cdot},D\big)=\frac{1}{|D|}\sum_{x\in D}\mathrm{RR}(x)$ (6)
@@ 3.4 Highlight Explanation Techniques
$\small\phi^{\text{FA}}_{i}=f_{a}(x)-f_{a}(x\setminus\{x_{i}\}\cup\{\tilde{x}_{i}\}),$ (7)
@@
$\small\phi^{\text{IG}}_{i}=(x_{i}-x^{\prime}_{i})\cdot\frac{1}{m}\sum_{k=1}^{m}\frac{\partial f_{a}\!\bigl(x^{\prime}+\tfrac{k}{m}(x-x^{\prime})\bigr)}{\partial x_{i}}\,.$ (8)
@@
$\small\phi^{\text{ATTN}}_{i}=A^{(L)}_{h^{\star},\,\text{gen},\,i}$ (10)
@@ 3.5 Mechanistic Interpretability for Highlight Explanations
$\small r^{(l,h)}\;=\;\bigl[\operatorname{Attn}^{(l,h)}_{\text{gen}}\bigr]\,W^{(l,h)}_{O}$, (11)
@@
$\small\text{logit}^{(l,h)}(a)\;=\;\langle W_{a},\,r^{(l,h)}\rangle\;=\;\bigl(W_{U}r^{(l,h)}\bigr)[a]$ (12)
@@
$\small S^{(l,h)}_{\tau}=\text{logit}^{(l,h)}(\text{Ans}_{\tau})-\text{logit}^{(l,h)}(\text{Ans}_{\tau^{\prime}}),$ (13)
$\small S^{(l,h)}_{\tau^{\prime}}=-S^{(l,h)}_{\tau}$ (14)
@@
$\small(l^{\star},h^{\star})\in\arg\max_{l,h}S^{(l,h)}_{c}\;\text{for }D_{C},$ (15)
$\small(l^{\star},h^{\star})\in\arg\max_{l,h}S^{(l,h)}_{m}\;\text{for }D_{M},$ (16)
@@
$\small\phi^{\text{MechLight}}_{i}\;=\;A^{(l^{\star})}_{h^{\star},\,\text{gen},\,i}$, (17)
@@ A.1 Datasets Details
Dataset Ctx. type #Inst. Avg ctx len
World Capital Conf. 55,830 37.9
Irre. 55,830 37.9
DoubleConf. 55,830 75.9
Mixed 55,830 75.9
Counterfact Conf. 802 44.8
Irre. 802 44.8
DoubleConf. 802 89.5
Mixed 802 89.5
Fakepedia Conf. 5,348 704.5
Irre. 5,348 704.5
DoubleConf. 5,348 1408.8
Mixed 5,348 1408.9
ConflictQA Conf. 1,343 593.1
Irre. 1,343 454.1
DoubleConf. 1,343 1190.2
Mixed 1,343 1047.2
@@ A.2 Other Details for Explanation Evaluation
$\small I\bigl(Y;X^{(k)}_{s}\bigr)\;=\;H(Y)\;-\;H\!\bigl(Y\mid X^{(k)}_{s}\bigr)$. (18)
@@
$\small H(Y)=\begin{cases}0,&p\in\{0,1\},\\ -p\log p-(1-p)\log(1-p),&p\in(0,1).\end{cases}$ (19)
@@
$\small\widehat{H}\!\bigl(Y\mid X^{(k)}_{s}\bigr)\;=\;\frac{1}{n}\sum_{i=1}^{n}h_{b}(\hat{p}_{i}).$ (21)
@@
$\small\widehat{I}\bigl(Y;X^{(k)}_{s}\bigr)\;=\;H(Y)\;-\;\widehat{H}\!\bigl(Y\mid X^{(k)}_{s}\bigr)$. (22)
@@
$\small\textsc{NMutInf@k};=\;\frac{\widehat{I}\bigl(Y;X^{(k)}_{s}\bigr)}{H(Y)}\;\in\;[0,1].$ (23)
@@ A.5 Faithfulness Evaluation Implementation Details
$\small\mathrm{AOPC}_{\text{comp}}\;=\;\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\big[\ell(x)-\ell(x^{\mathrm{mask}\,k})\big],$ (24)
$\small\mathrm{AOPC}_{\text{suff}}\;=\;\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\big[\ell(x)-\ell(x^{\mathrm{keep}\,k})\big].$ (25)

### v2

@@ Abstract
Context utilisation, the ability of Language Models (LMs) to incorporate relevant information from the provided context when generating responses, remains largely opaque to users, who cannot determine whether models draw from parametric memory or provided context, nor identify which specific context pieces inform the response. Highlight explanations (HEs) offer a natural solution as they can point the exact context pieces and tokens that influenced model outputs. However, no existing work evaluates their effectiveness in accurately explaining context utilisation. We address this gap by introducing the first gold standard HE evaluation framework for context attribution, using controlled test cases with known ground-truth context usage, which avoids the limitations of existing indirect proxy evaluations. To demonstrate the framework’s broad applicability, we evaluate four HE methods – three established techniques and MechLight, a mechanistic interpretability approach we adapt for this task – across four context scenarios, four datasets, and five LMs. Overall, we find that MechLight performs best across all context scenarios. However, our findings reveal systematic failures in all methods: explanation accuracy degrades significantly with context length, and all methods exhibit strong positional biases in multi-document settings. Surprisingly, widely used gradient-based methods provide little value for understanding context usage. These results challenge HEs’ utility in retrieval-augmented generation, factual verification, and other applications. Our framework provides the foundation for developing accurate context attribution methods.
@@ 1 Introduction
Figure 1: Utility evaluation of two HEs in our framework under Conflicting and Double-Conflicting context setups. In the left-hand example, the model selects the answer from the given passage. Explainer 2 shows better utility than Explainer 1. In the right-hand example, the model selects the answer from passage 2. Explainer 1 shows better utility than Explainer 2.
Language models (LMs) are increasingly deployed in applications requiring integration of provided context with parametric knowledge – from retrieval-augmented generation and question answering to document analysis and fact-checking. However, a fundamental transparency gap remains: users cannot determine whether model outputs draw from provided context or internal memory  (Jin et al., 2024; Yu et al., 2023; Monea et al., 2024a), nor identify which specific context informed the response. Highlight explanations (HEs) address this need naturally by pinpointing portions of the context responsible for the generation. See Fig.1 for two examples of HEs with high/low accuracy. Although HEs have proven valuable for understanding model decisions across various tasks (Sun et al., 2025; Ray Choudhury et al., 2023; Atanasova et al., 2020), no existing work evaluates their effectiveness in accurately explaining context utilisation.
@@
Across five LMs and four commonly used context-usage datasets, we find that MechLight HEs perform best across all context scenarios. However, two systematic limitations persist across all HEs: (i) length sensitivity – HE accuracy degrades as context grows; and (ii) position biases under dual‑context inputs: FA/IG tend to favour later (near‑question) pieces, while ATTN/MechLight favour earlier pieces. Surprisingly, the widely used IG and ATTN exhibit poor accuracy in most context scenarios, rendering them useless in revealing the model’s context utilisation. These failures also underscore the urgent need for explanation techniques that maintain accuracy at scale and overcome positional biases in multi-document settings. Our framework provides the foundation for future development of accurate context usage explanation methods.
@@ 2.1 Studies of Context Usage
Unsupervised Context Usage Explanations. To attribute the model’s answer to the specific part of the context, SelfCite trains a classifier on pseudo‑citations generated by the LLM itself (Chuang et al., ); ContextCite scores each sentence by the drop in answer likelihood when it is masked (Cohen-Wang et al., 2024). They require extra supervision, expensive perturbations, and only explain at the sentence level.
@@ 3.3 Metrics
$\begin{split}\Delta Rank@k^{\mathrm{grp}}(T;A,B)&=Rank@k(T,D_{B})-Rank@k(T,D_{A})\end{split}$ (1)
@@
$\begin{split}\Delta Rank@k^{\mathrm{inst}}_{D_{C_{a}}}=\frac{1}{|{D}_{C_{a}}|}\sum_{x\in{D}_{C_{a}}}(Rank@k^{\mathrm{inst}}(c_{b},x)-Rank@k^{\mathrm{inst}}(c_{a},x))\end{split}$ (2)
@@
$MDL-Bits@k=L_{\mathrm{preq}}(Y\mid X^{(k)}_{s})$ (4)
@@
$\mathrm{RR}(x)=\nicefrac{{1}}{{\mathrm{rank}\!\big(\text{Ans}_{\cdot};x\big)}}$ (5)
$MRR\big(T{=}\text{Ans}_{\cdot},D\big)=\frac{1}{|D|}\sum_{x\in D}\mathrm{RR}(x)$ (6)
@@ 3.4 Highlight Explanation Techniques
$\phi^{\text{FA}}_{i}=f_{a}(x)-f_{a}(x\setminus\{x_{i}\}\cup\{\tilde{x}_{i}\}),$ (7)
@@
$\phi^{\text{IG}}_{i}=(x_{i}-x^{\prime}_{i})\cdot\frac{1}{m}\sum_{k=1}^{m}\frac{\partial f_{a}\!\bigl(x^{\prime}+\tfrac{k}{m}(x-x^{\prime})\bigr)}{\partial x_{i}}\,.$ (8)
@@
$\phi^{\text{ATTN}}_{i}=A^{(L)}_{h^{\star},\,\text{gen},\,i}$ (10)
@@ 3.5 Mechanistic Interpretability for Highlight Explanations
$r^{(l,h)}\;=\;\bigl[\operatorname{Attn}^{(l,h)}_{\text{gen}}\bigr]\,W^{(l,h)}_{O}$, (11)
@@
$\text{logit}^{(l,h)}(a)\;=\;\langle W_{a},\,r^{(l,h)}\rangle\;=\;\bigl(W_{U}r^{(l,h)}\bigr)[a]$ (12)
@@
$S^{(l,h)}_{\tau}=\text{logit}^{(l,h)}(\text{Ans}_{\tau})-\text{logit}^{(l,h)}(\text{Ans}_{\tau^{\prime}}),$ (13)
$S^{(l,h)}_{\tau^{\prime}}=-S^{(l,h)}_{\tau}$ (14)
@@
$(l^{\star},h^{\star})\in\arg\max_{l,h}S^{(l,h)}_{c}\;\text{for }D_{C},$ (15)
$(l^{\star},h^{\star})\in\arg\max_{l,h}S^{(l,h)}_{m}\;\text{for }D_{M},$ (16)
@@
$\phi^{\text{MechLight}}_{i}\;=\;A^{(l^{\star})}_{h^{\star},\,\text{gen},\,i}$, (17)
@@ A.1 Datasets Details
Dataset Ctx. type #Inst. Avg ctx len World Capital Conf. 55,830 37.9 Irre. 55,830 37.9 DoubleConf. 55,830 75.9 Mixed 55,830 75.9 Counterfact Conf. 802 44.8 Irre. 802 44.8 DoubleConf. 802 89.5 Mixed 802 89.5 Fakepedia Conf. 5,348 704.5 Irre. 5,348 704.5 DoubleConf. 5,348 1408.8 Mixed 5,348 1408.9 ConflictQA Conf. 1,343 593.1 Irre. 1,343 454.1 DoubleConf. 1,343 1190.2 Mixed 1,343 1047.2
@@ A.2 Other Details for Explanation Evaluation
$I\bigl(Y;X^{(k)}_{s}\bigr)\;=\;H(Y)\;-\;H\!\bigl(Y\mid X^{(k)}_{s}\bigr)$. (18)
@@
$H(Y)=\begin{cases}0,&p\in\{0,1\},\\ -p\log p-(1-p)\log(1-p),&p\in(0,1).\end{cases}$ (19)
@@
$\widehat{H}\!\bigl(Y\mid X^{(k)}_{s}\bigr)\;=\;\frac{1}{n}\sum_{i=1}^{n}h_{b}(\hat{p}_{i}).$ (21)
@@
$\widehat{I}\bigl(Y;X^{(k)}_{s}\bigr)\;=\;H(Y)\;-\;\widehat{H}\!\bigl(Y\mid X^{(k)}_{s}\bigr)$. (22)
@@
$\textsc{NMutInf@k};=\;\frac{\widehat{I}\bigl(Y;X^{(k)}_{s}\bigr)}{H(Y)}\;\in\;[0,1].$ (23)
@@ A.5 Faithfulness Evaluation Implementation Details
$\mathrm{AOPC}_{\text{comp}}\;=\;\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\big[\ell(x)-\ell(x^{\mathrm{mask}\,k})\big],$ (24)
$\mathrm{AOPC}_{\text{suff}}\;=\;\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\big[\ell(x)-\ell(x^{\mathrm{keep}\,k})\big].$ (25)